ABCD is a parallelogram
AB = 2x + 6
BC = 8
CD = x + 8
Find the perimeter of ABCD

Found 2 solutions by fcabanski, amalm06:
Answer by fcabanski(1391)

Opposite sides of a parallelogram are equal. BC=8 so AD=8. AB = CD....solve for x and find AB and CD.


2x+6 = x+8 --->x=2. AB is 2(2)+6 =10 and CD is also 10 but plug in x to make sure 2+8=10.


The perimeter is 8+8+10+10 = 36

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)


Answer by amalm06(224)

Since opposite sides of a parallelogram are congruent, 2x+6=x+8
So that x=2
AB=DC=10
Perimeter=(2)(10)+(2)(8)=20+16=36